OpenStudy (anonymous):

Find the slope-intercept form of the line that passes through given the point and has the given slope. (1, 3) m = 3

Nnesha (nnesha):

y = mx + b this is a slope intercept form where m is slope which is always with x and b is y-intercept so replace m by 3 use that order pair where first number is x and 2nd one is y ( x ,y) and replace x and y by their values

Nnesha (nnesha):

( 1 , 3) what is x and y in this order pair ??

OpenStudy (anonymous):

X is 1 and Y is 3

Nnesha (nnesha):

yep right so replace x by 1 and y by 3 and then solve for b :-) y = 3x + b

OpenStudy (anonymous):

How do I solve for b???

Nnesha (nnesha):

okay what you get after substituting x and y values

Nnesha (nnesha):

did*

OpenStudy (anonymous):

3=3(1)+b

Nnesha (nnesha):

yep right multiply 3(1)

OpenStudy (anonymous):

3

Nnesha (nnesha):

3 times 1 =3 so 3 = 3 + b now you have to solve for b so that's how move 3 to the left side